# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v21

@@ 3.17.2 Speech-to-Text
On the converse, Nadungodage et al. [313] have done a series of work on Sinhala speech recognition with special notice given to Sinhala being a resource poor language. This project divides its focus on: continuity [314], active learning [315], and speaker adaptation [316]. A Sinhala speech recognition for voice dialing which is speaker independent was proposed by Amarasingha and Gamini [317] and on the other end, a Sinhala speech recognition methodology for interactive voice response systems, which are accessed through mobile phones was proposed by Manamperi et al. [318]. A Sinhala speech to Unicode text converter for the disaster relief domain was proposed by Prasangini and Nagahamulla [319]Priyadarshani [320] proposes a method for speaker dependant speech recognition based on their previous work on: dynamic time warping for recognizing isolated Sinhala words [321], genetic algorithms [322], and syllable segmentation method utilizing acoustic envelopes [323]. The method proposed by Gunasekara and Meegama [324] utilizes an HMM model for Sinhala speech-to-text. A Sinhala speech recognizer supporting bi-directional conversion between Unicode Sinhala and phonetic English was proposed by Punchimudiyanse and Meegama [325]. The work by Karunanayake et al. [326] transfer learns CNNs for transcribing free-form Sinhala and Tamil speech data sets for the purpose of classification. Dilshan [327] conducted a study for the specific use case of transcribing number sequences in continuous Sinhala speech.  Gamage et al. [328] explored the use of combinational acoustic models such as Deep Neural Network - Hidden Markov Model (DNN-HMM) [329] and Subspace Gaussian Mixture Model (SGMM) [329] in Sinhala speech recognition. In the follow-up work, Gamage et al. [330] extended that work with end-to-end Lattice-Free Maximum Mutual Information (e2e LF-MMI) model [331] which is claimed to be a viable solution for low resource language speech recognition by Carmantini et al. [332]. However, it was shown that the new model slightly underperforms compared to the state-of-the-art result.  Karunathilaka et al. [333] explore Sinhala speech recognition using deep learning models such as: pre-trained DNN, DNN, TDNN, TDNN+LSTM. The first half of the study by Arafath [334] dealt with recognizing Sinhala speech using LSTMs. Gamage et al. [310] included a Sinhala speech recognition module as one of the three modules present in their currency recognition system. Time-delay neural network architectures (including multistream CNN architecture) were used for acoustic modeling of Sinhala Automatic Speech Recognition (ASR) by Warusawithana et al. [335]. They have used the Kaldi speech recognition toolkit [336] for training the ASR models. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [337], proposed a Sinhala speech recognition system. The study by Azir et al. [338] attempts to identify number sequences spoken in Sinhala. TacoSi introduced by Arachchige and Weerasinghe [339] is based on Tacotron [340] and has been evaluated with 10 human evaluators to determine its text-to-speech quality. Nanayakkara [341] used DeepSpeech102102102https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition. Gunarathne et al. [342] used an earlier version of the CMUSphinx toolkit103103103https://cmusphinx.github.io/ to transcribe Sinhala speech to text. The later work by Akesh and Meegama [343] also used CMUSphinx toolkit on features extracted from Mel-frequency cepstral coefficients (MFCC) [344, 345] to automatically generate Sinhala subtitles from Speech.
@@ 3.17.5 Speech classification
The Sinhala speech classification system proposed by Buddhika et al. [209] does so without converting the speech-to-text. However, they report that this approach only works for specific domains with well-defined limited vocabularies. The work by Dinushika et al. [351] uses automatic speech recognition of Sinhala for speech command classification. Extending that, Kavmini et al. [352] presented a Sinhala speech command classification system which can be used for downstream applications. The work by Welarathna et al. [353] used CNNs to classify emotions (sad, disgust, surprise, neutral, happy, calm, fear, and angry) of Sinhala speech by Autistic children. The voice assistant system created by Senarathne et al. [354] is capable of handling Sinhala voice commands.
@@ 3.18.1 Printed Text
The XTREME-UP104104104https://github.com/google-research/xtreme-up data set created by Ruder et al. [130] contains a Sinhala data set for the OCR task. The data was obtained from book transcriptions. An attempt for Sinhala OCR system has been taken by Rajapakse et al. [357] before any other work has been done on the topic. Much later, a linear symmetry-based approach was proposed by Premaratne and Bigun [358, 359]. They then used hidden Markov model-based optimization on the recognized Sinhala script [360]. Similarly, Hewavitharana et al. [361] used hidden Markov models for off-line Sinhala character recognition. Herath et al. [112], Herath and Medagoda [113] developed a prepossessing engine based on a template matching for printed Sinhala OCR. Statistical approaches with histogram projections for Sinhala character recognition is proposed by Hewavitharana and Kodikara [362], by Ajward et al. [363], and by Madushanka et al. [364].  Karunanayaka et al. [365] also did off-line Sinhala character recognition with a use case for postal city name recognition. A separate group had attempted Sinhala OCR [366] mainly involving the nearest-neighbor method [367, 368]. A study by Ediriweera [369] uses dictionaries to correct errors in Sinhala OCR. An early attempt for Sinhala OCR by Dias et al. [370] has been extended to be online and made available to use via desktops [371] and hand-held devices [372] with the ability to recognize handwriting. A simple neural network based approach for Sinhala OCR was utilized by Rimas et al. [373]. A fuzzy-based model for identifying printed Sinhala characters was proposed by Gunarathna et al. [374]Premachandra et al. [375] proposes a simple back-propagation artificial neural network with hand crafted features for Sinhala character recognition. Another neural network with specialized feature extraction for Sinhala character recognition was proposed by Jayamaha and Naleer [376]. On the matter of neural networks and feature extraction, a feature selection process for Sinhala OCR was proposed by Kumara and Ragel [377].  Jayawickrama et al. [378] worked on Sinhala printed characters with special focus on handling diacritic vowels. However, they opted to refer to diacritic vowels as modifiers in their work. Gunawardhana and Ranathunga [379] proposed a limited approach to recognize Sinhala letters on Facebook images. A CNN-based methodology to improve printed Sinhala character OCR was proposed by Liyanage [380]. Printed Character Recognition (PCR) was used by Vasantharajan and Thayasivam [102] to create a large-scale Tamil-Sinhala-English parallel corpus. A meta-study on the effects of text genre, image resolution, and algorithmic complexity needed for Sinhala OCR from books and newspapers was conducted by Anuradha et al. [381].  Anuradha et al. [382] used Tesseract 3105105105https://tesseract-ocr.github.io/ [383] for Sinhala OCR. The later study by [384] improved the accuracy of Tesseract OCR engine on Sinhala from 53.22% to 86.16% for the data set they tested on.  Maduranga and Jayalal [385] used Artificial Neural Network (ANN) based on Universe of Discourse and Self Organization Map methods to recognize multi-style printed Sinhala characters. The study by De Silva et al. [386], even though ostensibly presented as translation research, is just invoking Google Cloud service for translation. Their novelty in Sinhala NLP lies in the facility provided to OCR the text from photos and documents. An OCR and Text-to-Speech system for Sinhala named Bhashitha was proposed by De Zoysa et al. [304]. A study on Sinhala text extraction from social media images (memes) was conducted by Samarajeewa and Ranathunga [387]. They specifically handled the character-touching issue. The study by Walawage and Ranathunga [388] attempted to devise a feature set to separately identify Sinhala and English text on social media images (memes). The study by de Silva and Liyanage [389] used Convolutional Spiking Neural Networks to extract Sinhala text from YouTube thumbnails.  Chanda et al. [390] proposed a Gaussian kernel SVM based method for word-wise Sinhala, Tamil, and English script identification. The work by Vasantharajan et al. [103] adapted the Tesseract engine to handle non-Unicode (legacy fonts) in pdf documents to create a Tamil-Sinhala-English parallel corpus.
@@ 3.18.2 Handwritten Text
Fernando et al. [391] claim to have created a database for handwriting recognition research in the Sinhala language and further claims that the data set is available at the National Science Foundation (NSF) of Sri Lanka. However, the paper provides no URLs and we were not able to find the data set on the NSF website either. The work by Karunanayaka et al. [392] is focused on noise reduction and skew correction of Sinhala handwritten words. A genetic algorithm-based approach for non-cursive Sinhala handwritten script recognition was proposed by Jayasekara and Udawatta [393]Nilaweera et al. [394] compare projection and wavelet-based techniques for recognizing handwritten Sinhala script.  Silva and Kariyawasam [395] worked on segmenting Sinhala handwritten characters with special focus on handling diacritic vowels. A comparative study of few available Sinhala handwriting recognition methods was done by Silva et al. [396]Silva et al. [397] uses contour tracing for isolated characters in handwritten Sinhala text. A Sinhala handwriting OCR system which utilizes zone-based feature extraction has been proposed by Dharmapala et al. [398]. The study by Walawage and Ranathunga [399] and its follow up study by Walawage [400] specifically focus on segmentation of overlapping and touching Sinhala handwritten characters. Silva and Jayasundere [401] focused on recognizing character modifiers in Sinhala handwriting. The similarly named studies by Mariyathas et al. [402] and Wasalthilake and Kartheeswaran [403], both utilize CNN to recognize Sinhala handwriting; as does the study by Weerasinghe [404]Ifhaam and Jayalal [405] used genetic algorithms to recognize Sinhala handwritten postal addresses for postal sorting. A segmentation-based approach that utilizes an n-gram model to recognize and validate Sinhala words written on touch screens was proposed by Mahesh and Priyankara [406]. They used a CNN classifier and were able to classify 19 different Sinhala characters. The study by Rowel et al. [407] frames their work as an E-Learning platform for hearing impaired children. However, their research does not contain any work done towards Sinhala sign language to be included in Section 3.23. What they do have is an OCR system that they claim to recognize letters and digits. Even there, we are only given an example of a recognised digit. Whether or not their system can recognize Sinhala letters is not explicitly discussed. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [337], proposed a Sinhala handwriting letter recognition system. The study by Withana and Rupasinghe [408] used Sinhala handwritten text classification to detect Dyslexia and Dysgraphia.
@@ 3.19 Translators
A meta-study on the viability of machine translators replacing English to Sinhala human translators was conducted by Dilshani and Senevirathna [418]. However, this study only involves 100 combined and complex English sentences translated to Sinhala by human translators as well as MT software. Given that reason and the fact that they seem to only used Google translate and Akura Sinhala dictionary app for comparison, the conclusions of this study may not be generalized. Another meta-study on the impact of pre-trained multilingual sequence-to-sequence models on low-resource language translation has been conducted by Lee et al. [419]; while they consider Sinhala as one of the examples, they do not go much into the specific impact due to the general nature of the paper. The meta-analysis by Ramadasa et al. [420] attempts to evaluate the goodness of the Google Sinhala-to-English translation by using the Google cloud API to translate Sinhala to English and then analysing the accuracy of the Sentiment Analysis task and the Named Entity Recognition task on the translated text. The meta-analysis by Das et al. [421] compared the results of translating English to 15 Indic languages including Sinhala using statistical translation methods. They used datasets from OPUS [119] for model building and utilized Flores-200 for fine-tuning. The NMT for Indic languages study conduced by Sheshadri et al. [422] discusses the Sinhala language translation in the abstract and conclusion but the paper itself focuses more on the languages spoken in India. Nevertheless, it puts Sinhala into a regional linguistic perspective. The meta-study by Bapna et al. [423] discusses the task of building clean, web-mined datasets for a number of languages including Sinhala for the task of machine translation. This discussion was continued by Jones et al. [424] who discussed the bilingual lexica (BILEXs) in the context of a number of languages including Sinhala. As mentioned in Section 3.1, the study by [93] raised questions on the quality of the existing Sinhala-English corpora. Sen et al. [425] on the other hand attempted to improve the quality of the Sinhala-English parallel corpora using fuzzy string matching where they tried to match the English translation of the given Sinhala sentence to the English sentence in the dataset pair. A study on the viability of using Google Translate for the legal domain English-Sinhala and Sinhala-English translation was conducted by [426]. They used human experts to extensively analyse the end result of the translation with many concrete examples of legal phrases that needed to be translated. The multilingual lexicon for low-resource machine translation dataset GATITOS introduced by Jones et al. [427] does not include Sinhala among their 170 languages set. However, they use Sinhala data from other sources and include several observations for Sinhala in this task. They observe that PanLex [428] and GATITOS improve results in the En-Si direction for smaller models (Transformer Big, 475M) but weaken the results in the Si-En direction. For bigger models (Transformer 1.6B), they do not observe an improvement or degradation of En-Si or Si-En direction.

### v22

@@ 3.17.2 Speech-to-Text
On the converse, Nadungodage et al. [327] have done a series of work on Sinhala speech recognition with special notice given to Sinhala being a resource poor language. This project divides its focus on: continuity [328], active learning [329], and speaker adaptation [330]. A Sinhala speech recognition for voice dialing which is speaker independent was proposed by Amarasingha and Gamini [331] and on the other end, a Sinhala speech recognition methodology for interactive voice response systems, which are accessed through mobile phones was proposed by Manamperi et al. [332]. A Sinhala speech to Unicode text converter for the disaster relief domain was proposed by Prasangini and Nagahamulla [333]Priyadarshani [334] proposes a method for speaker dependant speech recognition based on their previous work on: dynamic time warping for recognizing isolated Sinhala words [335], genetic algorithms [336], and syllable segmentation method utilizing acoustic envelopes [337]. The method proposed by Gunasekara and Meegama [338] utilizes an HMM model for Sinhala speech-to-text. A Sinhala speech recognizer supporting bi-directional conversion between Unicode Sinhala and phonetic English was proposed by Punchimudiyanse and Meegama [339]. The work by Karunanayake et al. [340] transfer learns CNNs for transcribing free-form Sinhala and Tamil speech data sets for the purpose of classification. Dilshan [341] conducted a study for the specific use case of transcribing number sequences in continuous Sinhala speech.  Gamage et al. [342] explored the use of combinational acoustic models such as Deep Neural Network - Hidden Markov Model (DNN-HMM) [343] and Subspace Gaussian Mixture Model (SGMM) [343] in Sinhala speech recognition. In the follow-up work, Gamage et al. [344] extended that work with end-to-end Lattice-Free Maximum Mutual Information (e2e LF-MMI) model [345] which is claimed to be a viable solution for low resource language speech recognition by Carmantini et al. [346]. However, it was shown that the new model slightly underperforms compared to the state-of-the-art result.  Karunathilaka et al. [347] explore Sinhala speech recognition using deep learning models such as: pre-trained DNN, DNN, TDNN, TDNN+LSTM. The first half of the study by Arafath [348] dealt with recognizing Sinhala speech using LSTMs. Gamage et al. [323] included a Sinhala speech recognition module as one of the three modules present in their currency recognition system. Time-delay neural network architectures (including multistream CNN architecture) were used for acoustic modeling of Sinhala Automatic Speech Recognition (ASR) by Warusawithana et al. [349]. They have used the Kaldi speech recognition toolkit [350] for training the ASR models. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [351], proposed a Sinhala speech recognition system. The study by Azir et al. [352] attempts to identify number sequences spoken in Sinhala. TacoSi introduced by Arachchige and Weerasinghe [353] is based on Tacotron [354] and has been evaluated with 10 human evaluators to determine its text-to-speech quality. Nanayakkara [355], Nanayakkara and Weerasinghe [356] used DeepSpeech110110110https://github.com/mozilla/DeepSpeech by Mozilla for Sinhala speech recognition. Gunarathne et al. [357] used an earlier version of the CMUSphinx toolkit111111111https://cmusphinx.github.io/ to transcribe Sinhala speech to text. The later work by Akesh and Meegama [358] also used CMUSphinx toolkit on features extracted from Mel-frequency cepstral coefficients (MFCC) [359, 360] to automatically generate Sinhala subtitles from Speech.
@@ 3.17.5 Speech classification
The Sinhala speech classification system proposed by Buddhika et al. [216] does so without converting the speech-to-text. However, they report that this approach only works for specific domains with well-defined limited vocabularies. The work by Dinushika et al. [366] uses automatic speech recognition of Sinhala for speech command classification. Extending that, Kavmini et al. [367] presented a Sinhala speech command classification system which can be used for downstream applications. The voice assistant system created by Senarathne et al. [368] is capable of handling Sinhala voice commands. The work by Welarathna et al. [369] used CNNs to classify emotions (sad, disgust, surprise, neutral, happy, calm, fear, and angry) of Sinhala speech by Autistic children. Sundarapperuma [370] created a speech emotion detection system for Sinhala.
@@ 3.18.1 Printed Text
The XTREME-UP112112112https://github.com/google-research/xtreme-up data set created by Ruder et al. [134] contains a Sinhala data set for the OCR task. The data was obtained from book transcriptions. An attempt for Sinhala OCR system has been taken by Rajapakse et al. [373] before any other work has been done on the topic. Much later, a linear symmetry-based approach was proposed by Premaratne and Bigun [374, 375]. They then used hidden Markov model-based optimization on the recognized Sinhala script [376]. Similarly, Hewavitharana et al. [377] used hidden Markov models for off-line Sinhala character recognition. Herath et al. [116], Herath and Medagoda [117] developed a prepossessing engine based on a template matching for printed Sinhala OCR. Statistical approaches with histogram projections for Sinhala character recognition is proposed by Hewavitharana and Kodikara [378], by Ajward et al. [379], and by Madushanka et al. [380].  Karunanayaka et al. [381] also did off-line Sinhala character recognition with a use case for postal city name recognition. A separate group had attempted Sinhala OCR [382] mainly involving the nearest-neighbor method [383, 384]. A study by Ediriweera [385] uses dictionaries to correct errors in Sinhala OCR. An early attempt for Sinhala OCR by Dias et al. [386] has been extended to be online and made available to use via desktops [387] and hand-held devices [388] with the ability to recognize handwriting. A simple neural network based approach for Sinhala OCR was utilized by Rimas et al. [389]. A fuzzy-based model for identifying printed Sinhala characters was proposed by Gunarathna et al. [390]Premachandra et al. [391] proposes a simple back-propagation artificial neural network with hand crafted features for Sinhala character recognition. Another neural network with specialized feature extraction for Sinhala character recognition was proposed by Jayamaha and Naleer [392]. On the matter of neural networks and feature extraction, a feature selection process for Sinhala OCR was proposed by Kumara and Ragel [393].  Jayawickrama et al. [394] worked on Sinhala printed characters with special focus on handling diacritic vowels. However, they opted to refer to diacritic vowels as modifiers in their work. Gunawardhana and Ranathunga [395] proposed a limited approach to recognize Sinhala letters on Facebook images. A CNN-based methodology to improve printed Sinhala character OCR was proposed by Liyanage [396]. Printed Character Recognition (PCR) was used by Vasantharajan and Thayasivam [106] to create a large-scale Tamil-Sinhala-English parallel corpus. A meta-study on the effects of text genre, image resolution, and algorithmic complexity needed for Sinhala OCR from books and newspapers was conducted by Anuradha et al. [397].  Anuradha et al. [398] used Tesseract 3113113113https://tesseract-ocr.github.io/ [399] for Sinhala OCR. The later study by [400] improved the accuracy of Tesseract OCR engine on Sinhala from 53.22% to 86.16% for the data set they tested on.  Maduranga and Jayalal [401] used Artificial Neural Network (ANN) based on Universe of Discourse and Self Organization Map methods to recognize multi-style printed Sinhala characters. The study by De Silva et al. [402], even though ostensibly presented as translation research, is just invoking Google Cloud service for translation. Their novelty in Sinhala NLP lies in the facility provided to OCR the text from photos and documents. An OCR and Text-to-Speech system for Sinhala named Bhashitha was proposed by De Zoysa et al. [317]. A study on Sinhala text extraction from social media images (memes) was conducted by Samarajeewa and Ranathunga [403]. They specifically handled the character-touching issue. The study by Walawage and Ranathunga [404] attempted to devise a feature set to separately identify Sinhala and English text on social media images (memes). The study by de Silva and Liyanage [405] used Convolutional Spiking Neural Networks to extract Sinhala text from YouTube thumbnails.  Chanda et al. [406] proposed a Gaussian kernel SVM based method for word-wise Sinhala, Tamil, and English script identification. The work by Vasantharajan et al. [107] adapted the Tesseract engine to handle non-Unicode (legacy fonts) in pdf documents to create a Tamil-Sinhala-English parallel corpus.
@@ 3.18.2 Handwritten Text
Fernando et al. [407] claim to have created a database for handwriting recognition research in the Sinhala language and further claims that the data set is available at the National Science Foundation (NSF) of Sri Lanka. However, the paper provides no URLs and we were not able to find the data set on the NSF website either. The work by Karunanayaka et al. [408] is focused on noise reduction and skew correction of Sinhala handwritten words. A genetic algorithm-based approach for non-cursive Sinhala handwritten script recognition was proposed by Jayasekara and Udawatta [409]Nilaweera et al. [410] compare projection and wavelet-based techniques for recognizing handwritten Sinhala script.  Silva and Kariyawasam [411] worked on segmenting Sinhala handwritten characters with special focus on handling diacritic vowels. A comparative study of few available Sinhala handwriting recognition methods was done by Silva et al. [412]Silva et al. [413] uses contour tracing for isolated characters in handwritten Sinhala text. A Sinhala handwriting OCR system which utilizes zone-based feature extraction has been proposed by Dharmapala et al. [414]. The study by Walawage and Ranathunga [415] and its follow up study by Walawage [416] specifically focus on segmentation of overlapping and touching Sinhala handwritten characters. Silva and Jayasundere [417] focused on recognizing character modifiers in Sinhala handwriting. The similarly named studies by Mariyathas et al. [418] and Wasalthilake and Kartheeswaran [419], both utilize CNN to recognize Sinhala handwriting; as does the study by Weerasinghe [420]Ifhaam and Jayalal [421] used genetic algorithms to recognize Sinhala handwritten postal addresses for postal sorting. A segmentation-based approach that utilizes an n-gram model to recognize and validate Sinhala words written on touch screens was proposed by Mahesh and Priyankara [422]. They used a CNN classifier and were able to classify 19 different Sinhala characters. The study by Rowel et al. [423] frames their work as an E-Learning platform for hearing impaired children. However, their research does not contain any work done towards Sinhala sign language to be included in Section 3.23. What they do have is an OCR system that they claim to recognize letters and digits. Even there, we are only given an example of a recognised digit. Whether or not their system can recognize Sinhala letters is not explicitly discussed. As part of their child [sic] cognitive ability assessment model, Kahawanugoda et al. [351], proposed a Sinhala handwriting letter recognition system. The study by Withana and Rupasinghe [424] used Sinhala handwritten text classification to detect Dyslexia and Dysgraphia. The study by eka [425] has used CNN to recognize Sinhala handwritten text.
@@ 3.19 Translators
A meta-study on the viability of machine translators replacing English to Sinhala human translators was conducted by Dilshani and Senevirathna [435]. However, this study only involves 100 combined and complex English sentences translated to Sinhala by human translators as well as MT software. Given that reason and the fact that they seem to only used Google translate and Akura Sinhala dictionary app for comparison, the conclusions of this study may not be generalized. Another meta-study on the impact of pre-trained multilingual sequence-to-sequence models on low-resource language translation has been conducted by Lee et al. [436]; while they consider Sinhala as one of the examples, they do not go much into the specific impact due to the general nature of the paper. The meta-analysis by Ramadasa et al. [437] attempts to evaluate the goodness of the Google Sinhala-to-English translation by using the Google cloud API to translate Sinhala to English and then analysing the accuracy of the Sentiment Analysis task and the Named Entity Recognition task on the translated text. The meta-analysis by Das et al. [438] compared the results of translating English to 15 Indic languages including Sinhala using statistical translation methods. They used datasets from OPUS [123] for model building and utilized Flores-200 for fine-tuning. The NMT for Indic languages study conducted by Sheshadri et al. [439] discusses the Sinhala language translation in the abstract and conclusion but the paper itself focuses more on the languages spoken in India. Nevertheless, it puts Sinhala into a regional linguistic perspective. The meta-study by Bapna et al. [440] discusses the task of building clean, web-mined datasets for a number of languages including Sinhala for the task of machine translation. This discussion was continued by Jones et al. [441] who discussed the bilingual lexica (BILEXs) in the context of a number of languages including Sinhala.
As mentioned in Section 3.1, the study by [96] raised questions on the quality of the existing Sinhala-English corpora. In a follow-up study by Ranathunga et al. [147], it was pointed out that using an automated ranking based on a similarity measure on web-mined corpora and using the resultant top samples can yield better translation models for English-Sinhala and Sinhala-Tamil. These models sometimes were better performing than a model trained on the full dataset. They further showed that using human labour to clean web-mined parallel corpora only gives marginal benefits over automated ranking and filtering. Thus they concluded that using expensive human labour for this task might not be efficient. Sen et al. [442] on the other hand attempted to improve the quality of the Sinhala-English parallel corpora using fuzzy string matching where they tried to match the English translation of the given Sinhala sentence to the English sentence in the dataset pair.
A study on the viability of using Google Translate for the legal domain English-Sinhala and Sinhala-English translation was conducted by [443]. They used human experts to extensively analyse the end result of the translation with many concrete examples of legal phrases that needed to be translated. The multilingual lexicon for low-resource machine translation dataset GATITOS introduced by Jones et al. [444] does not include Sinhala among their 170 languages set. However, they use Sinhala data from other sources and include several observations for Sinhala in this task. They observe that PanLex [445] and GATITOS improve results in the En-Si direction for smaller models (Transformer Big, 475M) but weaken the results in the Si-En direction. For bigger models (Transformer 1.6B), they do not observe an improvement or degradation of En-Si or Si-En direction.